OpenStudy (anonymous):

Isabelle made fifty-one thousand dollars last year. Randall made twenty thousand dollars plus thirty percent commission on ninety thousand dollars in sales. Who had a higher income last year? Explain.

OpenStudy (experimentx):

Randall 20k+30%of 90k = 20+0.3*90 = 47k